OpenStudy (anonymous):

The acceleration of a particle at time t moving along the x-axis is given by: a =4e^2t. At the instant when t=0, the particleis at the point x=2 moving with velocity v=-2. The position of the particle at t=1/2 is?

OpenStudy (mr.math):

You have to integrate \(a=4e^{2t}\) twice in order to get the position function.The first integral represents the velocity of the particle. Keep in mind that you have the point \(v(0)=-2\), which will enable you to find the constant of the first integral.

OpenStudy (mr.math):

Now, integrate the value you found again to get the the function that represents the position of the article. Use the given point \(x(0)=2\) to find the integration constant.

OpenStudy (anonymous):

\[\int\limits 4e ^{2t} dt\]

OpenStudy (anonymous):

that sould be v(t)

OpenStudy (mr.math):

Exactly!

OpenStudy (anonymous):

and then I would have to find C

OpenStudy (anonymous):

to do that i plug in -2

OpenStudy (mr.math):

You plug t=0 and v=-2.

OpenStudy (agreene):

\[\vec{r''}(t)=\vec{a}(t)\] \[\vec{r'}(t)=\vec{v}(t)\] \[\therefore \int\limits_{t_i}^{t_f} \vec{a}(t)dt=\vec v(t)\implies\int\limits_{t_i}^{t_f}\vec v(t) dt=\vec r(t)\] You're given the inital conditions: \[\vec{r}(0)=2m; \vec{v}(0)=-2ms^{-1}\] and \[\vec{a}(t)=4e^2t\] Just plug and chug.

OpenStudy (anonymous):

Thank You guys